the mean of 6numbers is 16. with the removal of a number the mean of remaining numbers is 17. the number removed is:

Grade:10

1 Answers

Ashutosh Sharma
8 years ago
initial mean=16
observations=6
new mean=17
observations=5
subtracted number= 16*6-17*5=96-85=11